Four pears and five apples cost $4.65. Three pears and four apples cost $3.60. Find the cost of a pear and the cost of an apple
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

4P+5A=4.65
3P+4A=3.60
============
16P+20A=18.60, multiply by 4
-15P-20A=-18.00 multiply by -5
P=$0.60
2.40+5A=4.65
5A=2.25
A=$0.45
-----------------
12P+15A=13.95
-12P-16A=-14.40. Multiply each by a number so that one of the terms will cancel out. Here, the P will
Add
-A=-0.45
Apples are $0.45
Pears are 3P+1.80=3.60
3P=1.80
P=$0.60
Apples are 45 cents each and pears are 60 cents each.

Answer by ikleyn(53937)

.
Four pears and five apples cost $4.65. Three pears and four apples cost $3.60. Find the cost of a pear and the cost of an apple
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

4p + 5a = 465,   (1)
3p + 4a = 360.   (2)

Solve it by the Elimination method. Multiply the equation (1) by 4 (both sides) and the equation (2) by 5, 
then distract the second from the first. You will get

16p - 15p = 465*4 - 360*5,

p = 60.

One pear costs 60 cents.

From this point, complete the solution on your own.